OpenStudy (miss.rose):

A line has a slope of –3/2 and has a y-intercept of 3. What is the x-intercept of the line?

OpenStudy (studygurl14):

First set up the equation y=mx + b m = slope b = y-intercept

OpenStudy (studygurl14):

To find the x-intercept, plug in 0 for y and solve for x

OpenStudy (miss.rose):

How do you solve?

OpenStudy (mrnood):

y=mx+b for x intercept y= 0 so 0 = mx+b you know m and b so put those values into the above and you can solve x

OpenStudy (studygurl14):

slope = -3/2 y -intercept = 3 y = mx + b y = (-3/2)x + 3 Plug in 0 for y and solve for x 0 = (-3/2)x + 3

OpenStudy (miss.rose):

x=2?

OpenStudy (studygurl14):

Perfect